The flight from Seattle to New York with tail wind takes five hours 15 minutes and the return flight takes six hours 35 minutes. since the flight time includes time to taxi from and to the gate we should subtract 45 minutes from each time to account for this . not shown, I was able to find out the plane was scheduled to fly a 2423 mile flight path. assuming the plane and the wind speeds are constant during both flights estimate the wind speed , to 1 decimal place
Answer by ankor@dixie-net.com(22740)

The flight from Seattle to New York with tail wind takes five hours 15 minutes and the return flight takes six hours 35 minutes. since the flight time includes time to taxi from and to the gate we should subtract 45 minutes from each time to account for this . not shown, I was able to find out the plane was scheduled to fly a 2423 mile flight path. assuming the plane and the wind speeds are constant during both flights estimate the wind speed , to 1 decimal place
:
Find the actual flying times each way
5:15 - :45 = 4:30 or 4.5 hrs
6:35 - :45 = 5:50 or 5.83 hrs
:
let s = plane speed in still air
let w = wind speed
Write a distance equation each way
4.50(s+w) = 2423
5.83(s-w) = 2423
-------------------
Divide the 1st equation by 4.5, divide the 2nd equation by 5.83
s + w = 538.44
s - w = 415.61
-------------------subtraction eliminates s, find w
0 + 2w = 122.83
w = 122.83/2
w ~ 61.4 mph is the speed of the wind